# Joint Transmit and Jamming Power Optimization for Secrecy in Energy Harvesting Networks: A Reinforcement Learning Approach

### v1

@@ IV-B Optimal Joint Power Allocation (OJPA)
In this section, we present an optimal approach called the optimal joint power allocation (OJPA) scheme for transmiter and jammer. Since the transitions to state $s^{(k+1)}$ depend solely on the current state $s^{(k)}$ and the current action $a^{(k)}$, our system follows the Markov property. Therefore, the proposed problem outlined in (12) is an online sequential decision-making problem with finite action, state spaces, and a bounded and consistent immediate reward function. We use MDP framework to obtain the solution where the goal is to make optimal decisions at each decision epoch to maximize the expected total reward [12] [26].
@@
Based on the geometric distribution of the lifetime of the network $K$, (16) is equivalent to the expected total discounted reward of infinite-horizon MDP [12, Proposition 5.3.1]
@@
$d^{*}(s)=\arg\max_{a\in U(s)}\bigg{\{}R(s,a)+\Gamma\sum_{s^{\prime}\in\cal{S}}\mathbb{P}(s^{\prime}|s,a)V(s^{\prime})\bigg{\}}.$ (19)
The pseudo-code for PI algorithm is given in Algorithm 1, which describes steps to finding optimal stationary deterministic policy $d^{*}(s)$ for each $s\in\mathcal{S}$ and storing these policies in a look-up table.
We refer to the phase of populating the look-up table in Algorithm 1 is known as the planning phase. The look-up table can then be used at each TS to allocate the power. The planning phase is further divided into two phases, i.e., policy evaluation and policy improvement as shown in Algorithm 1. The policy evaluation phase computes $V(s)$ for a given policy $d(s)$ by updating $V(s)$ iteratively from its initial value given in line number one for each $s\in\mathcal{S}$ using the Bellman equation in line number six until it converges in line number nine. Next, the policy improvement phase finds a better policy $d(s)$ than the given policy $\hat{d}(s)$ for each state $s\in\mathcal{S}$. The policy $d(s)$ is obtained by choosing the action $a\in U(s)$ that maximizes $V(s)$ corresponding to $\hat{d}(s)$ in line number thirteen. If the policy is unstable ($\hat{d}(s)\neq d(s)$), we repeat the policy evaluation phase again with a better policy obtained in line number thirteen. The algorithm continues iterating between the policy evaluation and policy improvement phase until the optimal policy is found, when, $\hat{d}(s)=d(s)$.
@@
Input: sets of state, action, state transition probability, and reward;
@@
5:         $v\leftarrow V(s)$
6:         $V(s)\leftarrow[R(s,a)+\Gamma\underset{s^{{}^{\prime}}\in\mathcal{S}}{\sum}\mathbb{P}(s^{{}^{\prime}}\mid s,a)V(s^{{}^{\prime}})]$
7:         $\Delta\leftarrow\max(\Delta,|v-V(s)|)$
@@
10:policy-stable $\leftarrow$ true
@@
12:     $\hat{d}(s)\leftarrow d(s)$
13:     $d(s)=\arg\underset{a\in U(s)}{\max}[R(s,a)+\Gamma\underset{s^{{}^{\prime}}\in\mathcal{S}}{\sum}\mathbb{P}(s^{{}^{\prime}}\mid s,a)V(s^{{}^{\prime}})]$
@@
15:         policy-stable $\leftarrow$ false
@@
Next, we refer to the subsequent phase of power allocation as the transmission phase. During this phase, the power allocation or action is obtained at each TS by directly fetching the power allocation values corresponding to the states from the look-up table populated by Algorithm 1. The pseudo-code of the transmission phase is described in Algorithm 2. In Algorithm 2, we iterate over the decision epochs to obtain the current state $s^{(k)}$ by generating the channel states and battery states first. Then, for each decision epoch, the optimal action $a^{(k)}$ is chosen based on the current state $s^{(k)}$ in line number seven of Algorithm 2 from the look-up table created in the planning phase. This provides the joint optimal $P^{(k)}_{\textrm{S}}$ and $P^{(k)}_{\textrm{D}}$ for transmission and jamming, respectively, in the $k$th TS. Using these power values, the cumulative reward is obtained and battery states are updated.
@@
9:     Calculate the total expected discounted reward $C^{(k)}_{\textrm{S}}$ for state $s^{(k)}$.
@@ IV-C1 Reduced State Joint Power Allocation (RSJPA)
In RSJPA, the algorithm chooses a smaller subset $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ by taking states randomly from $\mathcal{S}$ and creates a look-up table for $\mathcal{S}^{\prime}$ using Algorithm 1 [10]. In the transmission phase, the power allocation is carried out from the look-up table for the states in $\mathcal{S}^{\prime}$ and for the remaining states $\mathcal{S}\setminus\mathcal{S}^{\prime}$, the GA in (20), explained in the next section, is applied. Although this approach results in a degraded performance than that of the OJPA algorithm, it dramatically decreases the computational complexity of the planning phase to obtain the look-up table. For example, with $N_{S}/2$ states in $\mathcal{S}^{\prime}$, we can reduce computation complexity by $\frac{N_{A}^{N_{S}/2}}{2}$ times as compared to that of the OJPA algorithm with $N_{S}$ states in $\mathcal{S}$.
@@ IV-C3 Naive Algorithm (NA)
The NA algorithm also does not require the planning phase. In the transmission phase, it fully utilizes the energy stored in the battery at node S and node D for transmission and jamming, respectively, in each TS [27] [28], i.e., the transmit and jamming power in the $k$th TS are $P^{(k)}_{\textrm{S}}=\frac{B^{(k)}_{\textrm{S}}}{T_{x}}$ and $P^{(k)}_{\textrm{D}}=\frac{B^{(k)}_{\textrm{D}}}{T_{x}}$, respectively.
For the practical implementation of the joint power allocation algorithms OJPA and RSJPA, a central processing unit with all the system information may create the look-up table for power allocation. The central processing unit may then direct S and D to configure their respective powers at each TS by referring to the look-up table corresponding to the system states.
@@ V-A Individual Transmit Power Allocation (ITPA)
In this case, we only optimize $P_{\textrm{S}}^{(k)}$ when $P_{\textrm{D}}^{(k)}=P_{\textrm{D}}$ for each TS. To this end, we modify the MDP accordingly. The state and action set of the system in the $k$th TS can now be represented as $s^{(k)}=(G_{\textrm{SD}}^{(k)},G_{\textrm{SE}}^{(k)},G_{\textrm{DD}}^{(k)},G_{\textrm{DE}}^{(k)},B_{\textrm{S}}^{(k)})$ and $a^{(k)}\in U(s^{(k)})=\big{\{}P_{\textrm{S}}^{(k)}\mid 0\leq P_{\textrm{S}}^{(k)}\leq\frac{B_{\textrm{S}}^{(k)}}{T_{x}}\big{\}}$, respectively. The transition probabilities change based on $s^{(k)}$ and $a^{(k)}$ by following (14). With these changes, to obtain optimal solution, we apply Algorithm 1 and Algorithm 2 and to obtain sub-optimal solution, we apply NA, GA and RSJPA algorithms.
@@ V-B Individual Jamming Power Allocation (IJPA)
In this case, we only optimize $P_{\textrm{D}}^{(k)}$ when $P_{\textrm{S}}^{(k)}=P_{\textrm{S}}$ for each TS. Following changes in section V-A, the state and action set of the system in the $k$th TS can be represented as $s^{(k)}=(G_{\textrm{SD}}^{(k)},G_{\textrm{SE}}^{(k)},G_{\textrm{DD}}^{(k)},G_{\textrm{DE}}^{(k)},B_{\textrm{D}}^{(k)})$ and $a^{(k)}\in U(s^{(k)})=\big{\{}P_{\textrm{D}}^{(k)}\mid 0\leq P_{\textrm{D}}^{(k)}\leq\frac{B_{\textrm{D}}^{(k)}}{T_{x}}\big{\}}$, respectively. The transition probabilities will change based on $s^{(k)}$ and $a^{(k)}$ by following (14). With these changes, optimal and sub-optimal algorithms are applied as in section IV.
@@ VI Computational Complexity Analysis
In the worst case, Algorithm 1 of the OJPA algorithm may need to consider all possible policies. For each state, there are $N_{A}$ actions to choose from. As there are $N_{S}$ states, the total number of possible policies is $N_{A}^{N_{S}}$. As $N_{A}$ and $N_{S}$ grow, the number of possible policies grows exponentially. During each iteration of policy improvement, the algorithm typically eliminates several sub-optimal policies. As each iteration of policy improvement reduces the remaining policy space, on average, the worst-case computational complexity of the planning phase of the OJPA algorithm is $\mathcal{O}\big{(}\frac{N_{A}^{N_{S}}}{N_{S}}\big{)}$ [29]. In the transmission phase, power allocation at each TS is implemented by directly fetching power allocation values from the look-up table for that TS. As there are $K$ TSs, the transmission phase complexity is $K$.
The complexity of the RSJPA algorithm in the planning phase is $\mathcal{O}\big{(}\frac{N_{A}^{{N_{S}}/{2}}}{{N_{S}}/{2}}\big{)}$ as its planning phase is exactly similar to that of the OJPA algorithm with only half of the states ${N_{S}}/{2}$ [10]. When it comes to transmission phase complexity, the worst case possibility is that none of the states in the transmission phase belong to the look-up table due to the arbitrary selection of states for the preparation of the look-up table in the planning phase. That is why the worst case complexity in the transmission phase is $KN_{A}$ as for each TS, GA algorithm is implemented. On the contrary, the best case possibility is that all the states in the transmission phase belong to the look-up table. In this case, the complexity would be $K$ as in the OJPA algorithm. In the average case, states for the half of the TSs might belong to the look-up table but the states for the other half might not. In this case, the average complexity would be $\frac{K}{2}(1+N_{A})$.
The GA does not require a planning phase. Its transmission phase complexity is $KN_{A}$ as $N_{A}$ computations are required to identify the best action among $N_{A}$ actions that maximizes the current reward at each TS [10]. As in the GA, the NA also does not require a planning phase. As we just use the maximum stored energy in the batteries for transmission in each TS, the transmission phase complexity of the NA is $K$ [28].
@@
OJPA $\mathcal{O}\big{(}\frac{N_{A}^{N_{S}}}{N_{S}}\big{)}$ $K$
RSJPA $\mathcal{O}\big{(}\frac{N_{A}^{{N_{S}}/{2}}}{{N_{S}}/{2}}\big{)}$ Best case: $K$
Average case: $\frac{K}{2}(1+N_{A})$
Worst case: $KN_{A}$
GA -- $KN_{A}$
NA -- $K$
@@ VII Results and Discussions
Stopping criteria for policy evaluation loop $\epsilon$ $0.07$
Length of TS $T_{S}$ 10 ms
Transmission time $T_{x}$ $5$ ms
One energy unit $2.5$ $\mu$J
Harvested energy $E_{H}$ $2$ energy units
Battery capacity of node S $B_{\textrm{S}}^{\textrm{max}}$ $5$ energy units
Battery capacity of node D $B_{\textrm{D}}^{\textrm{max}}$ $5$ energy units
Probability of harvesting $E_{H}$ units of energy at S $p$ $0.5,0.8$
Probability of harvesting $E_{H}$ units of energy at D $q$ $0.5,0.8$
Set of transmit and jamming power $\mathcal{P}$$\{0,0.5,1,2\}$ mW
Set of transmit and jamming energy $\mathcal{E_{U}}$$\{0,1,2,4\}$ energy unit
@@
$\eta_{E}=\mathbb{E}\left[\mathbb{E}_{K}\left[\frac{\sum_{k=0}^{K-1}C_{\textrm{S}}^{(k)}T_{x}}{\sum_{k=0}^{K-1}(P_{\textrm{S}}^{(k)}+P_{\textrm{D}}^{(k)})T_{x}}\right]\right].$ (21)
@@
In Fig. 2(a) we compare the expected total discounted reward versus $\Gamma$ for the algorithms OJPA, RSJPA, GA, and NA when the probability of EH improves from $p=q=0.5$ to $p=q=0.8$. The corresponding energy efficiency plot is shown in Fig. 2(b). The motivations behind plotting Fig. 2(a) and Fig. 2(b) are three-fold. First is to check whether a higher value of $\Gamma$ leads to a larger expected total discounted reward and energy efficiency or not. As $1/(1-\Gamma)$ is the average lifetime of the network, a higher value of $\Gamma$ indicates a longer lifetime. Thus, a higher value of $\Gamma$ should lead to a better performance. The second is to assess the relative performance of algorithms, OJPA, RSJPA, GA, and NA. Third is to find how EH probability $p$ and $q$ affect performance.
@@
It can be observed from Figs. 2(a) and 2(b) that the higher value of $\Gamma$ leads to better expected total discounted reward and energy efficiency. From Fig. 2(a), it can be observed that the OJPA algorithm performs the best and the NA performs the worst as expected. The OJPA algorithm takes into account the long-term performance of the system and that is why its performance is the best. The NA performs the worst because it does not consider both immediate and future rewards. Instead, it simply utilizes all the stored energy in the battery at every TS. The GA outperforms the NA as it prioritizes maximizing immediate reward. The performance of the RSJPA algorithm is close to the OJPA algorithm being better than the GA which is near the OJPA algorithm as it adopts a hybrid approach between the OJPA algorithm and the GA. We observe from Fig. 2(a) that as $p$ and $q$ improve, all the algorithms tend to perform better in terms of the expected total discounted reward. In contrast, we notice from Fig. 2(b) that the energy efficiency for all the algorithms is better when the probability of EH decreases except for the NA. This observation prompts us to plot Figs. 3(a) and 3(b) to closely study how the expected total discounted reward and energy efficiency vary with the probability of EH and the reasoning for the same. A common observation from both Figs. 2(a) and 2(b) is that the performance gap between the OJPA algorithm and other algorithms is greater when the probability of EH decreases. This suggests that the OJPA algorithm is more beneficial at a lower probability of EH.
(a) Expected total discounted reward versus $p$ and $q$ for joint power allocation algorithms where $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units and $\Gamma=0.9$.
(b) Energy efficiency versus $p$ and $q$ for joint power allocation algorithms where $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units and $\Gamma=0.9$.
@@
Figure 3: Expected total discounted reward and energy efficiency versus EH probability.

### v2

@@ IV-B Optimal Joint Power Allocation (OJPA)
In this section, we present an optimal approach called the optimal joint power allocation (OJPA) scheme for transmiter and jammer. Since the transitions to state $s^{(k+1)}$ depend solely on the current state $s^{(k)}$ and the current action $a^{(k)}$, our system follows the Markov property. Therefore, the proposed problem outlined in (12) is an online sequential decision-making problem with finite action, state spaces, and a bounded and consistent immediate reward function. We use MDP framework to obtain the solution where the goal is to make optimal decisions at each decision epoch to maximize the expected total reward [12] [30]  [^1] Our proposed RL approach can be extended to multi-antenna systems, however, the curse of dimensionality, due to increased number of states, is a challenge. Leveraging deep RL techniques offers promising solutions to effectively address these difficulties..
@@
Based on the geometric distribution of the lifetime of the network $K$, (16) is equivalent to the expected total discounted reward of an infinite-horizon MDP [12, Proposition 5.3.1]
@@
$d^{*}(s)=\mathop{\mathrm{argmax}}_{a\in U(s)}\left\{R(s,a)+\Gamma\sum_{s^{\prime}\in\mathcal{S}}\mathbb{P}(s^{\prime}\mid s,a)V(s^{\prime})\right\}.$ (19)
The pseudo-code for the PI algorithm is given in Algorithm 1, which describes steps to finding optimal stationary deterministic policy $d^{*}(s)$ for each $s\in\mathcal{S}$ and storing these policies in a look-up table.
We refer to the phase of populating the look-up table in Algorithm 1 is known as the planning phase. The look-up table can then be used at each TS to allocate the power. The planning phase is further divided into two phases, i.e., policy evaluation and policy improvement, as shown in Algorithm 1. The policy evaluation phase computes $V(s)$ for a given policy $d(s)$ by updating $V(s)$ iteratively from its initial value given in line number one for each $s\in\mathcal{S}$ using the Bellman equation in line number six until it converges in line number nine. Next, the policy improvement phase finds a better policy $d(s)$ than the given policy $\hat{d}(s)$ for each state $s\in\mathcal{S}$. The policy $d(s)$ is obtained by choosing the action $a\in U(s)$ that maximizes $V(s)$ corresponding to $\hat{d}(s)$ in line number thirteen. If the policy is unstable ($\hat{d}(s)\neq d(s)$), we repeat the policy evaluation phase again with a better policy obtained in line number thirteen. The algorithm continues iterating between the policy evaluation and policy improvement phase until the optimal policy is found, when, $\hat{d}(s)=d(s)$.
@@
Input: Set of states, actions, state transition probability, and reward;
@@
5:   $v=V(s)$
6:   $V(s)=[R(s,a)+\Gamma\underset{s^{{}^{\prime}}\in\mathcal{S}}{\sum}\mathbb{P}(s^{{}^{\prime}}\mid s,a)V(s^{{}^{\prime}})]$
7:   $\Delta=\max(\Delta,|v-V(s)|)$
@@
10:policy-stable $=$ true
@@
12:  $\hat{d}(s)=d(s)$
13:  $d(s)=\mathop{\mathrm{argmax}}\limits_{a\in U(s)}\left[R(s,a)+\Gamma\sum_{s^{{}^{\prime}}\in\mathcal{S}}\mathbb{P}(s^{{}^{\prime}}\mid s,a)V(s^{{}^{\prime}})\right]$
@@
15:   policy-stable $=$ false
@@
Next, we refer to the subsequent phase of power allocation as the transmission phase. During this phase, the power allocation or action is obtained at each TS by directly fetching the power allocation values corresponding to the states from the look-up table populated by Algorithm 1. The pseudo-code of the transmission phase is described in Algorithm 2.  [^2] A CPU in the network which has the look-up table for power allocation from Algorithm 1 along with global channel states and battery states at the beginning of each TS, can execute Algorithm 2 as the decision-maker of the power allocation policies. In Algorithm 2, we iterate over the decision epochs to obtain the current state $s^{(k)}$ by generating the channel states and battery states first. Then, for each decision epoch, the optimal action $a^{(k)}$ is chosen based on the current state $s^{(k)}$ in line number seven of Algorithm 2 from the look-up table created in the planning phase. This provides the joint optimal $P^{(k)}_{\textrm{S}}$ and $P^{(k)}_{\textrm{D}}$ for transmission and jamming, respectively, in the $k$th TS. Using these power values, the cumulative reward is obtained and battery states are updated.
@@
9:  Calculate the total expected discounted reward $\Gamma^{k}C^{(k)}_{\textrm{S}}$ for state $s^{(k)}$.
@@ IV-C1 Reduced State Joint Power Allocation (RSJPA)
To develop a reduced complexity algorithm, we propose the RSJPA algorithm, which combines the OJPA algorithm outlined in Algorithm 1 with the GA described in (20). The RSJPA algorithm chooses a smaller subset $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ by taking states randomly from $\mathcal{S}$ and creates a look-up table for $\mathcal{S}^{\prime}$ using Algorithm 1 in the planning phase [10]. The reason for randomly selecting a state is that each state in the system has an equal probability. Therefore, choosing any subset of states will not affect the overall performance. In the transmission phase, the power allocation is carried out from the look-up table for the states in $\mathcal{S}^{\prime}$ and for the remaining states $\mathcal{S}\setminus\mathcal{S}^{\prime}$, the GA is applied. This approach strikes a balance between performance and complexity in comparison to the OJPA algorithm. By increasing the number of states included in the subset $\mathcal{S}^{\prime}$, we can significantly increase performance; however, this improvement comes with a rise in complexity. This trade-off highlights the importance of carefully considering the state selection to optimize outcomes effectively. For example, with $N_{S}/2$ states in $\mathcal{S}^{\prime}$, we can reduce computation complexity by $\frac{N_{A}^{N_{S}/2}}{2}$ times as compared to that of the OJPA algorithm with $N_{S}$ states in $\mathcal{S}$.
@@ IV-C3 Naive Algorithm (NA)
The NA algorithm also does not require the planning phase. In the transmission phase, it fully utilizes the energy stored in the battery at node S and node D for transmission and jamming, respectively, in each TS [31] [32], i.e., the transmit and jamming power in the $k$th TS are $P^{(k)}_{\textrm{S}}=\frac{B^{(k)}_{\textrm{S}}}{T_{s}}$ and $P^{(k)}_{\textrm{D}}=\frac{B^{(k)}_{\textrm{D}}}{T_{s}}$, respectively.
Implementing the algorithms proposed in this paper requires global channel state information, which includes the channel state information related to the eavesdropper. In certain scenarios, it may be possible to acquire the eavesdropper’s channel state information when the eavesdropper is an active node in the network and its transmissions can be monitored [33, 34]. For instance, in networks where nodes serve dual roles, acting as legitimate receivers for some transmissions while functioning as eavesdroppers for others, the channel state information of the eavesdroppers may be obtained. Another example is found in networks where confidential information is intended solely for a specific user, treating all other nodes as potential eavesdroppers, as seen in military communications. In such cases, any data transmission from the eavesdropper to the source and destination could enable the estimation of the eavesdropper’s channel state information by leveraging the reciprocal characteristics of the wireless channel. Furthermore, in the secure communication literature, it is a common assumption that channel state information related to eavesdroppers is available [28, 21, 22, 23, 24, 25, 26, 27, 33, 34, 35, 36, 37].
For the practical implementation of the joint power allocation algorithms OJPA and RSJPA, a CPU with access to all the system information, fed back from both the source and destination, can generate a look-up table for power allocation. The CPU can then instruct the source and destination to configure their respective power levels at each TS based on the look-up table corresponding to the system states.
@@ V-A Individual Transmit Power Allocation (ITPA)
In this case, we only optimize $P_{\textrm{S}}^{(k)}$ when $P_{\textrm{D}}^{(k)}=P_{\textrm{D}}$ for each TS. To this end, we modify the MDP accordingly. The state and action set of the system in the $k$th TS can now be represented as $s^{(k)}=(G_{\textrm{SD}}^{(k)},G_{\textrm{SE}}^{(k)},G_{\textrm{DD}}^{(k)},G_{\textrm{DE}}^{(k)},B_{\textrm{S}}^{(k)})$ and $a^{(k)}\in U(s^{(k)})=\big{\{}P_{\textrm{S}}^{(k)}\mid 0\leq P_{\textrm{S}}^{(k)}\leq\frac{B_{\textrm{S}}^{(k)}}{T_{s}}\big{\}}$, respectively. The transition probabilities change based on $s^{(k)}$ and $a^{(k)}$ by following (14). With these changes, to obtain optimal solution, we apply Algorithm 1 and Algorithm 2 and to obtain sub-optimal solution, we apply NA, GA and RSJPA algorithms.
@@ V-B Individual Jamming Power Allocation (IJPA)
In this case, we only optimize $P_{\textrm{D}}^{(k)}$ when $P_{\textrm{S}}^{(k)}=P_{\textrm{S}}$ for each TS. Following changes in section V-A, the state and action set of the system in the $k$th TS can be represented as $s^{(k)}=(G_{\textrm{SD}}^{(k)},G_{\textrm{SE}}^{(k)},G_{\textrm{DD}}^{(k)},G_{\textrm{DE}}^{(k)},B_{\textrm{D}}^{(k)})$ and $a^{(k)}\in U(s^{(k)})=\big{\{}P_{\textrm{D}}^{(k)}\mid 0\leq P_{\textrm{D}}^{(k)}\leq\frac{B_{\textrm{D}}^{(k)}}{T_{s}}\big{\}}$, respectively. The transition probabilities will change based on $s^{(k)}$ and $a^{(k)}$ by following (14). With these changes, optimal and sub-optimal algorithms are applied as in section IV.
@@ VI Computational Complexity Analysis
In the worst case, Algorithm 1 of the OJPA algorithm may need to consider all possible policies. For each state, there are $N_{A}$ actions to choose from. As there are $N_{S}$ states, the total number of possible policies is $N_{A}^{N_{S}}$. As $N_{A}$ and $N_{S}$ grow, the number of possible policies grows exponentially. During each iteration of policy improvement, the algorithm typically eliminates several sub-optimal policies. As each iteration of policy improvement reduces the remaining policy space, on average, the worst-case computational complexity of the planning phase of the OJPA algorithm is $\mathcal{O}\big{(}\frac{N_{A}^{N_{S}}}{N_{S}}\big{)}$ [38]. In the transmission phase, power allocation at each TS is implemented by directly fetching power allocation values from the look-up table for that TS. As there are $K$ TSs, the transmission phase complexity is $\mathcal{O}(K)$.
The complexity of the RSJPA algorithm in the planning phase with $\rho\%$ of $N_{S}$ states in $\mathcal{S}^{\prime}$, is given by $\mathcal{O}\big{(}\frac{N_{A}^{{\rho N_{S}}/{100}}}{{\rho N_{S}}/{100}}\big{)}$. This is because the planning phase of the RSJPA is implemented by executing the OJPA algorithm with ${\rho N_{S}}/{100}$ states [10]. When it comes to transmission phase complexity, the worst case possibility is that none of the states in the transmission phase belong to the look-up table due to the arbitrary selection of states for the preparation of the look-up table in the planning phase. That is why the worst case complexity in the transmission phase is $\mathcal{O}(KN_{A})$ as for each TS, GA algorithm is implemented. On the contrary, the best case possibility is that all the states in the transmission phase belong to the look-up table. In this case, the complexity would be $\mathcal{O}(K)$ as in the OJPA algorithm. In the average case, states for the $\rho\%$ of the TSs might belong to the look-up table but the states for the remaining $(100-\rho)\%$ TSs might not. In this case, the average complexity would be $\mathcal{O}(\frac{K}{100}(\rho+(100-\rho)N_{A}))$.
The GA does not require a planning phase. Its transmission phase complexity is $\mathcal{O}(KN_{A})$ as $N_{A}$ computations are required to identify the best action among $N_{A}$ actions that maximize the current reward at each TS [10]. As in the GA, the NA also does not require a planning phase. As we just use the maximum stored energy in the batteries for transmission in each TS, the transmission phase complexity of the NA is $\mathcal{O}(K)$ [32].
@@
OJPA $\mathcal{O}\big{(}\frac{N_{A}^{N_{S}}}{N_{S}}\big{)}$ $\mathcal{O}(K)$
RSJPA $\mathcal{O}\big{(}\frac{N_{A}^{{\rho N_{S}}/{100}}}{{\rho N_{S}}/{100}}\big{)}$ Best case: $\mathcal{O}(K)$
Average case: $\mathcal{O}(\frac{K}{100}(\rho+(100-\rho)N_{A}))$
Worst case: $\mathcal{O}(KN_{A})$
GA -- $\mathcal{O}(KN_{A})$
NA -- $\mathcal{O}(K)$
@@ VII Results and Discussions
Stopping criteria for policy evaluation loop $\epsilon$ 0.070.070.07
Duration of TS $T_{s}$ 5ms
One energy unit 2.52.52.5 $\mu$J
Harvested energy {$E_{\textrm{S}}$, $E_{\textrm{D}}$} {111, 222}, {222, 111} energy units
Battery capacity of node S $B_{\textrm{S}}^{\textrm{max}}$ 555 energy units
Battery capacity of node D $B_{\textrm{D}}^{\textrm{max}}$ 555 energy units
Probability of harvesting $E_{\textrm{S}}$ units of energy at S $p$ 0.5,0.80.5,0.80.5 , 0.8
Probability of harvesting $E_{\textrm{D}}$ units of energy at D $q$ 0.5,0.80.5,0.80.5 , 0.8
[…]
@@
$\eta_{E}=\mathbb{E}\left[\mathbb{E}_{K}\left[\frac{\sum_{k=0}^{K-1}C_{\textrm{S}}^{(k)}T_{s}}{\sum_{k=0}^{K-1}(P_{\textrm{S}}^{(k)}+P_{\textrm{D}}^{(k)})T_{s}}\right]\right].$ (21)
@@
(c) Expected total discounted reward versus $\Gamma$ for joint power allocation algorithms for different $p$ and $q$ with $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
(d) Energy efficiency versus $\Gamma$ for joint power allocation algorithms for different $p$ and $q$ with $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
@@
In Fig. 2(a) we compare the expected total discounted reward versus $\Gamma$ for the algorithms OJPA, RSJPA, GA, and NA when the probability of EH improves from $p=q=0.5$ to $p=q=0.8$ while harvested energy $E_{\textrm{S}}=E_{\textrm{D}}=2$. The corresponding energy efficiency plot is shown in Fig. 2(b). The motivations behind plotting Fig. 2(a) and Fig. 2(b) are four-fold. First is to check whether a higher value of $\Gamma$ leads to a larger expected total discounted reward and energy efficiency or not. As $1/(1-\Gamma)$ is the average lifetime of the network, a higher value of $\Gamma$ indicates a longer lifetime. Thus, a higher value of $\Gamma$ should lead to a better performance. The second is to assess the relative performance of algorithms, OJPA, RSJPA, GA, and NA. The third is to find how EH probability $p$ and $q$ affect performance when they are equal. Lastly, is to study the impact of varying the number of states in $\mathcal{S}^{\prime}$ on the performance of RSJPA.
@@
It can be observed from Figs. 2(a) - 2(d) that the higher value of $\Gamma$ leads to better expected total discounted reward and energy efficiency. From Fig. 2(a), it can be observed that the OJPA algorithm performs the best and the NA performs the worst as expected. The OJPA algorithm takes into account the long-term performance of the system and that is why its performance is the best. The NA performs the worst because it does not consider both immediate and future rewards. Instead, it simply utilizes all the stored energy in the battery at every TS. The GA outperforms the NA as it prioritizes maximizing immediate reward. The performance of the RSJPA algorithm is close to the OJPA algorithm being better than the GA which is near the OJPA algorithm as it adopts a hybrid approach between the OJPA algorithm and the GA.
We observe from Fig. 2(a) that as $p$ and $q$ improve, all the algorithms tend to perform better in terms of the expected total discounted reward. In contrast, we notice from Fig. 2(b) that the energy efficiency for all the algorithms is better when the probability of EH decreases except for the NA. This observation prompts us to plot Fig. 4 to closely study how the expected total discounted reward and energy efficiency vary with the probability of EH and the reasoning for the same. A common observation from both Figs. 2(a) and 2(b) is that the performance gap between the OJPA algorithm and other algorithms is greater when the probability of EH decreases. This suggests that the OJPA algorithm is more beneficial at a lower probability of EH.
[…]
(a) When $p$ improves from 0.50.50.5 to 0.80.80.8 with $q=0.5$ and $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
(b) When $q$ improves from 0.50.50.5 to 0.80.80.8 with $p=0.5$ and $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
(c) When $E_{\textrm{S}}$ improves from 111 to 222 with $E_{\textrm{D}}=2$ and $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
(d) When $E_{\textrm{D}}$ improves from 111 to 222 with $E_{\textrm{S}}=2$ and $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
@@
Figure 3: Expected total discounted reward versus discount factor $\Gamma$ with unequal probability of EH and harvested energy units at S and D.
Figs. 3(a) and 3(b) depict the same performance metric as of Fig. 2(a) when $p$ and $q$ are unequal keeping $E_{\textrm{S}}=E_{\textrm{D}}=2$, and Figs. 3(c) and 3(d) depict the same considering $E_{\textrm{S}}$ and $E_{\textrm{D}}$ to be unequal, when $p=q=0.8$. In Fig. 3(a), $p$ increases from 0.50.50.5 to 0.80.80.8 when $q=0.5$, whereas in Fig. 3(b), $q$ increases from 0.50.50.5 to 0.80.80.8 when $p=0.5$. In both figures, increasing EH probability either at S or D improves performance. However, the performance improvement is greater when EH probability improves at S. The observation in Figs. 3(c) and 3(d) with the change in harvested energy is similar, that is, when harvested energy increases from 1 to 2 energy units at S rather than at D, the performance improvement is more. We can conclude from Fig. 3 that the improvement of the probability of EH and the harvested energy at both S and D is beneficial for the system; however, the improvement of these metrics at S has a greater impact on the improvement of the performance of the system.
(a) When $q=0.5$, $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units and $\Gamma=0.9$.
(b) When $q=0.5$, $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units and $\Gamma=0.9$.
(c) When $p=0.5$, $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units and $\Gamma=0.9$.
(d) When $p=0.5$, $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units and $\Gamma=0.9$.
Figure 4: Expected total discounted reward and energy efficiency versus EH probability $p$ and $q$.